You deposit $3,000 in a retirement account today at 5.5% interest. How much more money will you have if you leave the money invested for forty-five years rather than forty years?

Select one: A. $7,839.74 B. $7,714.91 C. $7,858.19 D. $7,799.08 E. $7,846.52
Answer #1

Answer- Deposit Amount = $ 3000 Future value = Present Value(1 + R)n

Future value Of $1 at End of year 45 = $1(1+5.5%)45 = 11.12655 $

Future value Of $1 at End of year 40 = $1(1+5.5%)40 = 8.513309 $

Hence Difference Between Future Values for Year 45 - Year 40 = 11.12655 - 8.513309 = $2.613245

Total Difference = $ 3000 * $2.613245 = $7839.74

Hence Correct Answer is A. $7839.74

Year Future Values of $1 @5.5%
0 1
1 1.055
2 1.113025
3 1.174241375
4 1.238824651
5 1.306960006
6 1.378842807
7 1.454679161
8 1.534686515
9 1.619094273
10 1.708144458
11 1.802092404
12 1.901207486
13 2.005773897
14 2.116091462
15 2.232476492
16 2.355262699
17 2.484802148
18 2.621466266
19 2.765646911
20 2.917757491
21 3.078234153
22 3.247537031
23 3.426151568
24 3.614589904
25 3.813392349
26 4.023128928
27 4.244401019
28 4.477843075
29 4.7241244
30 4.983951288
31 5.258068609
32 5.547262383
33 5.852361814
34 6.174241714
35 6.513825008
36 6.872085383
37 7.250050079
38 7.648802834
39 8.06948699
40 8.513309
41 8.981540757
42 9.475525498
43 9.996679401
44 10.54649677
45 11.12655